# Conceptualization, Methodology, Writing – review & editing, Supervision

### v3

@@ 5 Results and Discussion
This section is organized into four subsections. First, the model performance evaluation based on crash severity classification (KABCO scale) is presented. Second, key contributing factors influencing pedestrian crashes near bus stops are described. Third, the specific impacts of these contributing factors on severe (KAB) pedestrian crashes are analyzed and discussed. Finally, the fourth subsection outlines the systematic identification and prioritization of hazardous bus stops using the developed framework.
@@ 5.2.5 Nearest intersection
Previous studies have primarily examined either intersections in general or only signalized intersections, consistently associating bus stops near these locations with increased pedestrian crash risks (Harwood et al., 2008; Kuşkapan et al., 2022; Walgren, 1998). Even at signalized intersections, bus stops within 1,000 feet have been linked to higher pedestrian crash risks compared to other spots (Geedipally, 2021; Quistberg et al., 2015b). This study extends prior research by simultaneously analyzing both signalized and non-signalized intersections. The results indicate that signalized intersections near bus stops are associated with a lower likelihood of pedestrian crashes, as reflected by the negative coefficient (-0.329). This suggests that traffic signals improve pedestrian safety by managing vehicle movements and reducing conflicts. These findings align with Samani and Amador-Jimenez (2023), who reported that stop signs at non-signalized intersections can also enhance safety by slowing vehicles and structuring traffic flow. The apparent divergence between this study’s findings and previous research may be due to differences in intersection characteristics, pedestrian compliance behaviors, and study locations. While intersections generally introduce higher pedestrian exposure, well-designed signalized intersections may mitigate risks more effectively than previously assumed. This underscores the importance of enhancing traffic control measures at non-signalized intersections to further improve pedestrian safety near bus stops. Stronger controls can help ensure more predictable interactions between pedestrians and vehicles, particularly in high-traffic areas, thereby reducing crash risks.
@@ 5.4 Hazard bus stops/corridors identification
However, the FB approach remains superior to both EB and Sim-EB methods, as it fully accounts for parameter uncertainty within a hierarchical Bayesian framework rather than relying solely on prior information from reference sites. By generating posterior samples without requiring closed-form solutions, FB provides greater accuracy in parameter estimation and predictions, particularly in complex models. Given these advantages, this study employs the FB method to ensure a comprehensive and precise identification of hazardous bus stops, establishing a robust framework for safety improvements. Expected crash frequencies were estimated, and hazardous bus stops were identified using the Potential for Safety Improvement (PSI) approach. PSI quantifies the difference between expected and predicted crashes, providing a structured measure of safety risk (Hauer et al., 2002). The expected crash frequency at each bus stop was calculated following Eqn 22, and for a more detailed understanding, readers are directed to Khodadadi et al., 2022. Notably, the PSI method has yet to be applied using the RPNB-L model, presenting an opportunity for future research to explore this innovative approach.
[…]

### v4

@@ 5 Results and Discussion
This section is divided into four subsections. First, it presents the model performance evaluation based on the KABCO crash severity scale. Second, it examines the key factors associated with pedestrian crashes near bus stops. Third, it explores the specific effects of these factors on severe crashes (KAB). Finally, the fourth subsection outlines the systematic identification and prioritization of high-risk bus stops using the proposed framework.
@@ 5.2.5 Nearest intersection
Intersections are widely recognized as critical zones for pedestrian risk due to the convergence of multimodal traffic streams and complex crossing behaviors. Numerous studies (Harwood et al., 2008; Kuşkapan et al., 2022; Walgren, 1998; Geedipally, 2021; Quistberg et al., 2015b) have reported elevated pedestrian crash frequencies at or near bus stops located within 1000 feet of signalized intersections. However, the scope of these investigations has primarily focused on signalized intersections, often overlooking unsignalized intersections, which may present heightened risk due to the lack of formal traffic control mechanisms. To address this limitation, the present study examined the relationship between pedestrian crash risk and intersection type, specifically comparing signalized and unsignalized intersections. The RPNB-L model results (Table 2) indicate a negative and statistically significant association between bus stops located near signalized intersections and pedestrian crashes (–0.329). This suggests that relative to unsignalized locations, signalized intersections may offer a protective effect by reducing crash likelihood. These findings align with those of Samani and Amador-Jimenez (2023), who observed that even modest infrastructure elements such as pedestrian signals and stop signs at non-signalized intersections can enhance driver awareness and pedestrian visibility. Signalized intersections likely reduce pedestrian risk by facilitating more orderly vehicle movement, minimizing turning conflicts, and providing designated pedestrian crossing phases.
To further validate these findings and explore potential confounding effects, a stratified Mantel–Haenszel (MH) analysis (Santner and Duffy, 2012) was conducted. This method provides an adjusted estimate of the association between crash occurrence and intersection type by accounting for key contextual variables across multiple strata. These strata were defined based on traffic volume (AADT), bus stop proximity to intersections, presence of marked crosswalks, and land use type. The MH estimator was computed using stratum-specific $2\times 2$ contingency tables, where each table cross-tabulated pedestrian crash occurrence (yes/no) with intersection type (signalized/unsignalized). The pooled Mantel–Haenszel odds ratio (MHOR) is given by:
$\hat{OR}_{MH}=\frac{\sum_{s=1}^{S}\frac{a_{s}d_{s}}{n_{s}}}{\sum_{s=1}^{S}\frac{b_{s}c_{s}}{n_{s}}}$ (22)
where $a_{s}$, $b_{s}$, $c_{s}$, and $d_{s}$ represent the cell counts of the $s$-th stratum, and $n_{s}$ is the total number of observations within that stratum.
Results from the MH analysis corroborated the model-based findings. Signalized intersections were consistently associated with significantly lower crash risk across all stratified analyses. For example, when stratified by AADT with different ranges, the MHOR was 0.284, corresponding to a relative risk reduction of approximately 41% (RR $\approx$ 0.59). Similarly, stratification by bus stop proximity on the far side yielded an MHOR of 0.313 (RR $\approx$ 0.62); by marked crosswalk presence, 0.272 (RR $\approx$ 0.56). All odds ratios were statistically significant at the $p<0.001$ level. These consistent results across multiple strata highlight the presence of traffic control mechanisms, such as signal phases, pedestrian indications, and regulated turning movements, may contribute to reducing pedestrian exposure to conflicts and enhancing overall safety. Moreover, these findings help reconcile discrepancies in previous literature by illustrating how aggregated or unstratified analyses may obscure the protective influence of signalization.
@@ 5.4 Hazard bus stops/corridors identification
However, the FB approach remains superior to both EB and Sim-EB methods, as it fully accounts for parameter uncertainty within a hierarchical Bayesian framework rather than relying solely on prior information from reference sites. By generating posterior samples without requiring closed-form solutions, FB provides greater accuracy in parameter estimation and predictions, particularly in complex models. Given these advantages, this study employs the FB method to ensure a comprehensive and precise identification of hazardous bus stops, establishing a robust framework for safety improvements. Expected crash frequencies were estimated, and hazardous bus stops were identified using the Potential for Safety Improvement (PSI) approach. PSI quantifies the difference between expected and predicted crashes, providing a structured measure of safety risk (Hauer et al., 2002). The expected crash frequency at each bus stop was calculated following Eqn 23, and for a more detailed understanding, readers are directed to Khodadadi et al., 2022. Notably, the PSI method has yet to be applied using the RPNB-L model, presenting an opportunity for future research to explore this innovative approach.
$p(y_{\text{exp}}\mid y_{\text{obs}})=\int_{\eta,\phi}NB\left(e^{-\eta},\phi\right)\left(\int_{b}p(\eta\mid b,y_{\text{obs}})\,\pi_{b}\,db\right)\,\pi_{\phi}\,d\eta\,d\phi$ (23)